TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Back to overview

YOU ARE VIEWING BlogEntryPoint8x2007x06x19x19x41

Powering Tests Sector 7-8 - 19.07.2007

Test Report from Tuesday 19.07.2007

  • RQD.A78:
    • Transfer function measurements with 60A converter completed OK
    • PCC Circuit Config tests completed
    • PIC2 - All steps completed OK except:
      • Circuit discharge from power converter: switches were not rearmed after PM data generation - Test to be repeated
      • Discharge request from PIC - to be done

  • RQF.A78:
    • Transfer function measurements with 60A converter completed OK
    • PCC Circuit Config tests completed
    • PIC2 - All steps completed OK up to Powering Failure test (included)

  • RQ4.L8 and RD2.L8:
    • Tests 1 and 2 proposed by MPP completed but without provoking an artificial quench at the end of the second. Detection threshold still at 20mV. Still no quench. Further research to be done.

  • RCO.A78B1 powered below 20A for QPS analysis. The circuit is still blocked till the new sofware is uploaded.

Test Plan for Wednesday 20.07.2007

  • Resolution of cable signal/interlock swap F/D

  • RQD.A78 (from the CCC):
    • Last two steps of PIC2
    • PLI1 (750A)
    • PLI2 (2000A)
    • PLI3 (6500A)

  • RQF.A78 (from the CCC):
    • Complete PIC2
    • First steps of PLI1,2,3 if possible

  • RB.A78
    • IST Energy extraction tests, PIC1, connection of the leads

  • Matching section:
    • Type analysis of the power converters
    • Analyse boil-off data and eventually repeat a more complete boil-off
    • Waiting for new proposals from MPP
    • 120A converter on Q5 type tests

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2007-06-20 - unknown
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login